# Reawakening knowledge: Anticipatory recovery from catastrophic interference via structured training

### v1

@@ 3.4 Anticipatory Recovery in Vision Models
For the causal image modeling task, we use pre-trained Image GPT models and the corresponding pixel tokenizers from Chen et al. (2020), and we use CIFAR-10 (Krizhevsky et al., 2009) as the fine-tuning dataset. Similar to the LLM experiments, we fine-tune our model on each sampled image for $M$ gradient steps, and repeat $E$ epochs with a fixed order of the images. The resulting loss curves are shown in Figure 6(a). The results show that the anticipatory recovery phenomenon also exists for sequential cyclic training of image modeling in addition to language modeling.
@@
For the image classification task, we use pre-trained vision transformer (ViT) (Dosovitskiy et al., 2020) and VGG-19 (Simonyan and Zisserman, 2015) models, and we use ImageNet (Deng et al., 2009) as the training dataset. For each experiment, we randomly sample 800 images from Imagenet and divide them into 25 batches of 32 images each. We train our model on each batch for $M$ gradient steps and repeat $E$ epochs with a fixed order of the batches. The resulting loss curves are plotted in Figure 6(b). Results show that both the transformer ViT and the convolutional VGG exhibit anticipatory recovery in cyclic training.
@@ 4 Understanding Cyclic Training Dynamics
(a) Gradients
(b) Amount of Recovery
(c) Weights
(d) Activations
Figure 8: Heat map visualizations for (a) cosine similarities between the gradient vectors of the attention layer in transformer block 12 of the model for each task; (b) loss recoveries for training on task $\bm{x}_{i}$ (y-axis) and evaluating on task $\bm{x}_{j}$ (x-axis); (c) cosine similarities between the flattened model weight vectors at each point in training; (d) cosine similarities between the last layer activations for document $\bm{x}_{1}$ at each point in training.
An important general question about the anticipatory recovery phenomenon is whether it is due to some causal mechanism relating the dynamics of model parameters to the training sequence, or whether it is more correlational in that adjacent tasks come to be represented more similarly by the model. We found initial evidence for the latter hypothesis in experiments locally reversing the task sequence (e.g., showing that $x_{t+1}$ primes $x_{t}$ nearly as much as vice versa). To further test this learned similarity hypothesis, we explore the relationships between tasks and across training history of the model’s loss gradients, weights, and activations. The results enable us to better understand the dynamics of cyclic training. Unless otherwise stated, all visualizations in this section use the 410M model and default hyperparameters.
@@ 4.1 Temporal Structure of Gradients
We do cyclic training for 4 epochs and compute the gradient of each document at the attention layer of transformer block 12 at the conclusion of training. In Figure 7(a), we plot the cosine similarities between these gradient vectors of each document. Results show that the gradients have mostly positive cosine similarities (except for the last document, on which the model has just been trained). We predicted the gradient similarities to depend only on $|i-j|$, peaking along the diagonal. Instead, to our surprise, the gradient similarities are highest near the center of the heat map. That is, the gradient similarity between documents $\bm{x}_{t-1}$ and $\bm{x}_{t}$ depends on where we are in the cycle. This result suggests an additional layer to the anticipatory recovery phenomenon: Recovery for document $\bm{x}_{t}$ is greatest from training on document $\bm{x}_{t-1}$, but the strength of the potential facilitation between $\bm{x}_{t-1}$ and $\bm{x}_{t}$ is actually greatest after we train for another $b$ documents (for some small number $b$, which ranges from roughly 10 to 15 in this particular case). We verify this by computing the pairwise recovery: we take the model checkpoint after 4 epochs of cyclic training, do $M$ gradient updates on document $\bm{x}_{i}$ and compute the difference in the loss of document $\bm{x}_{j}$ before and after these gradient updates, for each pair of documents $(\bm{x}_{i},\bm{x}_{j})$. We plot these pairwise loss recoveries in Figure 7(b). Results confirm that the amount of recovery on document $\bm{x}_{j}$ is highest when the model checkpoint is taken from roughly 10 to 15 documents before or after document $\bm{x}_{j}$ in cyclic training and then fine-tuned on a proximal document $\bm{x}_{i}$ in the sequence. The fact that this pairwise loss recovery matrix is roughly symmetric also suggests that the anticipatory recovery phenomenon approximately exhibits task symmetry: gradient updates on document $\bm{x}_{t}$ decrease the loss for document $\bm{x}_{t+k}$ for small integers $k$, and vice versa. We provide additional visualizations for $T=50,100,200$ in Appendix C and a more detailed description for this phenomenon.
Refer to caption Figure 9: Top three PCA components of last layer weights in the first three epochs.
Refer to caption (a) $f_{i}(\bm{w})=\bm{w}$ Refer to caption (b) $f_{i}(\bm{w})=\bm{y}_{i}-\bm{w}$
@@ 4.2 Temporal Structure of Model Weights
We explore the similarities of model weights along the optimization trajectory of cyclic training. We flatten and concatenate the weight vectors of the model checkpoint after we fine-tune on each document. However, the cosine similarities between these raw model weight vectors are all very close to 1 without obvious structure. This is mainly because of the proximity of model weights along the same optimization trajectory. Another reason is that the size of the weight vector is huge and dominated by a few entries, leading to numerical instability. To resolve these issues, we instead explore the structure of “weight residuals.” We compute the weight residuals by subtracting the average of weights in a window of length $T$ centered at the current document from the current weight, i.e.
$w_{\text{res}}(t)=w(t)-\frac{1}{T}\sum_{n=t-T/2}^{t+T/2}w(n).$ (2)
This removes the shared components along the optimization trajectory and allows us to focus on the model weight updates for each document. Figure 7(c) visualizes a heat map of the cosine similarity between each pair of weight residuals from the second epoch to the fourth epoch. The visualization shows a cyclic structure in the weight residuals, as equidistant bright stripes that align with the training epochs. Furthermore, each stripe spans several documents, suggesting the similarity of weight residuals for proximal documents.
In addition to computing the cosine similarities between model weights, we explore using Principle Component Analysis (PCA) to reduce the dimensionality of the weights. We compute the top three PCs of the flattened last-layer weight vector (the output word embedding layer) for the Pythia-1B model, and plot its trajectory in Figure 9. The plot exhibits a clear helical structure that gradually converges. We believe this is highly relevant to anticipatory recovery: right before revisiting a task, the projected model weights in the helix move closer to the point corresponding to the previous appearance of that task, leading to anticipatory recovery on the loss of that task. As we go on with cyclic training, the model also exhibits less forgetting and gradually converges to a solution that achieves low loss on all tasks.
It is important to note that the helical structure of the weight trajectory is not an obviously necessary consequence of the cyclical training. Cyclical training could be expected to yield a repeating pattern, but the facts that the tasks come to be organized in a circle that respects their ordering and that the trajectory goes through one full revolution per epoch (rather than some other arc length) are nontrivial and seem to be essential for anticipatory recovery.
@@ 4.4 Computational Toy Model
To further understand the essential mechanism that yields anticipatory recovery, we design a minimalist “toy” simulation experiment. In this toy simulation, each task (formerly, document) $i\in\{1,\cdots,T\}$ is described by a single data point, $\bm{x}_{1},\cdots,\bm{x}_{T}\in\mathbb{R}^{N}$. We assume a learnable linear embedding $\bm{P}\in\mathbb{R}^{M\times N}$ that projects each $\bm{x}_{i}$ into an $M$-dimensional embedding space. We also assume a learnable vector $w$ and task specific mappings $f_{i}$, where $f_{i}(w)$ is the target for task $i$ in the same embedding space. We require each $f_{i}$ to be invertible as a simplifying assumption.
We define the loss for task $i$ as
$\ell_{i}(\bm{P},\bm{w})=\frac{1}{2}\lVert\bm{P}\bm{x}_{i}-f_{i}(\bm{w})\rVert^{2}_{2}.$ (3)
Just as when training a deep net, we assume here that representation learning occurs slowly, and that one training step for task $i$ involves a single gradient update of $\bm{P}$ with step size $\alpha$:
@@
As in our main experiments, we sequentially optimize each $\ell_{i}$ as we iterate through the sequence of tasks. In each training step, we first update $\bm{P}$ and then solve for $\bm{w}$ given the update to $\bm{P}$. In essence, updating $\bm{P}$ approximately reduces the distance between $\bm{P}\bm{x}_{i+1}$ and $f_{i+1}(f_{i}^{-1}(\bm{P}\bm{x}_{i}))$. Assume each $f_{i}$ is Lipschitz, then this will entail bringing $f_{i+1}^{-1}(\bm{P}\bm{x}_{i+1})$ toward $f_{i}^{-1}(\bm{P}\bm{x}_{i})$. As a result of this optimization objective, the model will evolve along the optimization trajectory such that the $f_{i}^{-1}(\bm{P}\bm{x}_{i})$ for all tasks $i$ gradually form a circular pattern. This gives an intuitive explanation on the anticipatory recovery phenomenon, since updaing $\bm{w}$ according to equation 5 will also bring it closer to $f_{i+1}^{-1}(\bm{P}\bm{x}_{i+1})$, thus reducing the loss on task $i+1$ and exhibits anticipatory recovery.
@@
There are two potential reasons large overparameterized networks might produce the anticipatory recovery in a way analogous to the toy simulation. First, as the scale of the network grows, it is likely that the network can develop task-specific parameters that quickly adapt to and memorize the new input data, corresponding to Equation 5. And when the fast memorization is achieved, the gradient descent dynamics of the slow weights push the representations of the two adjacent tasks ($P\bm{x}_{i}$ and $P\bm{x}_{i+1}$) closer when $f$ is an identity function, according to Equation 4. In earlier LLM experiments, this effect can be seen in Figure 2, where larger models achieve significantly lower losses within a few gradient update steps during sequential learning. Second, larger networks have more learning capacity to map the features of two adjacent tasks closer. In our linear projection model, anticipatory recovery keeps growing over many epochs, whereas in LLM experiments, the anticipatory effect is already at the strongest within two or three epochs. Moreover, in the toy model all data points are randomly generated, which makes it easier to separate and map their representations according to a temporal structure than real-world data. In contrast, real-world data could require even more representation capacity since data points are noisy and correlated.
@@ 5 Related Work
Prior theoretical works have studied convergence rates, under various assumptions, for the training setup where the data points are shuffled only once and that order is reused for all epochs (Ahn et al., 2020; Gurbuzbalaban et al., 2019; Mishchenko et al., 2020; Safran and Shamir, 2020). On the empirical side, Xu et al. (2022) found that shuffling the data only once in the beginning can achieve a convergence rate comparable to shuffling every epoch. The training setup is equivalent to our cyclic training setup, but our research examines the loss on each task throughout the training cycle and discovers the anticipatory recovery effect. We also extend it to multiple gradient update steps on each data point.
Our research also relates to the more general topic of learning in structured environments. Jones et al. (2023) studied regression and classification tasks with multi-scale temporal structure in the environment characterized by $1/f$ dynamics. While the cyclic training setting that we study is a more simplified setup than that of Jones et al. (2023), we aim at unveiling more insights on applying standard SGD on over-parameterized networks. A potential direction for future work would be to study anticipatory recovery in regimes with richer, hierarchical sequence structure.
@@
Online learning deals with the setting where the tasks come from an online sequential stream. One of the simplest algorithms in online learning is follow-the-leader (Hannan, 1957), which stores all previous data from the stream and minimizes the total loss. It has strong performance guarantees but is computationally very expensive, and it also might not be feasible to store all the past data. Many subsequent works have developed cheaper algorithms under different assumptions (Zinkevich, 2003; Cesa-Bianchi and Lugosi, 2006; Shalev-Shwartz et al., 2012). Many recent works also explore the connection between online learning and meta-learning or continual learning (Denevi et al., 2019a; Finn et al., 2019; Denevi et al., 2019b; Javed and White, 2019; Fini et al., 2020; Ren et al., 2021; Wang et al., 2021).
The cyclic training setting that we explore in this research can be considered as a special case of the online learning setting where the data stream has a cyclic repetition structure. We employ multiple steps of online gradient descent (Biehl and Schwarze, 1995) on each document from the stream and study the training dynamics of over-parameterized neural networks.
LLM Emergent Capabilities.
Recent advancements in large-scale Transformer networks (Vaswani et al., 2017; Devlin et al., 2019) have demonstrated exceptional ability to model long sequence language data. Beyond basic language modeling and downstream task performance, these models have shown emergent behaviors (Wei et al., 2022a) that appear to manifest only beyond a certain model scale (Brown et al., 2020; Wei et al., 2022b; Ganguli et al., 2022; Srivastava et al., 2022).
Related to our research, recent studies reveal that LLMs possess remarkable memorization skills, enabling them to recall news sentences after just a few exposures (Biderman et al., 2023; Carlini et al., 2023; Orhan, 2023). However, the sequential learning dynamics behind such memorization have not been thoroughly examined. Our work comprehensively explore the sequential learning setting with cyclic task repetition and demonstrates task anticipation, a new emergent capability of large models.
@@
Traditionally, neural networks are trained for a single objective or task. However, when transitioning between tasks sequentially, they often experience “catastrophic interference” (McCloskey and Cohen, 1989), marked by a significant drop in performance on previously learned tasks. Continual learning (Van de Ven et al., 2020), a subset of machine learning, addresses a simplified setup where a model sequentially learns a set of tasks without revision. Numerous algorithms have been proposed to mitigate catastrophic forgetting, focusing on methods like parameter regularization (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018), data replay (Rebuffi et al., 2017; Rolnick et al., 2019; Chaudhry et al., 2019), knowledge distillation (Hinton et al., 2015; Li and Hoiem, 2017; Buzzega et al., 2020; Madaan et al., 2023), and architectural isolation and expansion (Yoon et al., 2018; Serra et al., 2018; Gurbuz and Dovrolis, 2022; Kang et al., 2022).
@@
More recently, there have been debates over the practicality of continual learning setups. Studies like Davidson and Mozer (2020) have shown that as networks learn more tasks, they improve in learning speed and reduce forgetting. In large models, studies suggest that pre-trained vision classifiers can undertake continual learning with ease, by either freezing or fine-tuning representations (Janson et al., 2022; Lee et al., 2023; Fini et al., 2022). In the language domain, research also suggests that LLMs exhibit emerging continual learning capabilities (Scialom et al., 2022; Ke et al., 2022).
Unlike prior literature on continual learning, our research uniquely focuses on sequential learning environments with cyclic repetition. Our work extends interleaved training (Mayo et al., 2023) to a larger number of tasks, specifically investigating the emergent anticipatory recovery phenomenon in cyclic training. This finding adds to the above literature by demonstrating a new mechanism by which large networks can avoid or recover from catastrophic interference.
6 Discussion

### v2

@@ 3.4 Anticipatory Recovery in Vision Models
Similar to the LLM experiments, we fine-tune a pre-trained Image GPT model chen2020generative on each sampled image from CIFAR-10 (krizhevsky2009learning, ) for $M$ gradient steps, and repeat $E$ epochs with a fixed order of the images. The resulting loss curves are shown in Figure 7(a). The results show that the anticipatory recovery phenomenon also exists for sequential cyclic training of image modeling in addition to language modeling.
@@
For each experiment, we randomly sample 800 images from Imagenet (deng2009imagenet, ) and divide them into 25 batches of 32 images each. We fine-tune a pre-trained vision transformer (ViT) (dosovitskiy2020image, ) and VGG-19 (simonyan2014very, ) model on each batch for $M$ gradient steps and repeat $E$ epochs with a fixed order of the batches. The resulting loss curves are plotted in Figure 7(b). Results show that both the transformer ViT and the convolutional VGG exhibit anticipatory recovery in cyclic training.
@@
3.5 Online Loss Evaluation
We compare the performance of training with fixed ordering and random shuffling of each epoch in the prequential evaluation setting with $T=100$ documents. Prequential evaluation (cai2021online, ) measures online performance by evaluating the model on the document it is about to be trained on, which is equivalent to evaluating the training loss of each batch. For the random shuffling condition, document 1 is always trained first in each epoch, and the other 99 documents are presented in a different random order in each epoch. In this experiment we set $C=512$, $M=10$, $E=5$ and the error bars are based on 10 different seeds.
The resulting loss curves are plotted in Figure 1(c), and the average loss throughout each run is summarized in Table 1. We exclude epoch 1 when computing the average loss since all documents are new to the model for both the cyclic training and random shuffling condition. We observe that training with fixed ordering is superior to random shuffling in the prequential evaluation setting across all 4 pre-trained Pythia models of different sizes, due to the structure in the data stream. The results suggest practical potential of structured training.
@@ 4 Understanding Cyclic Training Dynamics
Figure 8: Heat map visualizations for (a) cosine similarities between the gradient vectors of the attention layer in transformer block 12 of the model for each task; (b) loss recoveries for training on task $\bm{x}_{i}$ (y-axis) and evaluating on task $\bm{x}_{j}$ (x-axis); (c) cosine similarities between the flattened model weight residuals at each point in training; (d) cosine similarities between the last layer activations for document $\bm{x}_{1}$ at each point in training.
An important general question about anticipatory recovery is whether it is due to some causal mechanism relating the dynamics of model parameters to the training sequence, or whether it is more correlational in that adjacent tasks come to be represented more similarly by the model. We found initial evidence for the latter hypothesis in experiments locally reversing the task sequence (e.g., showing that $\bm{x}_{t+1}$ primes $\bm{x}_{t}$ nearly as much as vice versa). To further test this learned similarity hypothesis, we explore the relationships between tasks and the model’s loss gradients, weights, and activations across training history. The results enable us to better understand the dynamics of cyclic training. Unless otherwise stated, all visualizations in this section use the 410M model and default hyperparameters.
@@ 4.1 Temporal Structure of Gradients
We do cyclic training for 4 epochs and compute the gradient of each document at the attention layer of transformer block 12 at the conclusion of training. In Figure 8(a), we plot the cosine similarities between these gradient vectors of each document. Results show that the gradients have mostly positive cosine similarities (except for the last document, on which the model has just been trained). To our surprise, the gradient similarities are highest near the center of the heat map rather than peaking along the diagonal. That is, the gradient similarity between documents $\bm{x}_{t-1}$ and $\bm{x}_{t}$ depends on where we are in the cycle. This result suggests an additional layer to the anticipatory recovery phenomenon: Recovery for document $\bm{x}_{t}$ is greatest from training on document $\bm{x}_{t-1}$, but the strength of the potential facilitation between $\bm{x}_{t-1}$ and $\bm{x}_{t}$ is actually greatest after we train for another $b$ documents (for some small number $b$). We verify this by computing the pairwise recovery: we take the model checkpoint after 4 epochs of cyclic training, and then for each pair of documents $(\bm{x}_{i},\bm{x}_{j})$, we do $M$ gradient updates on $\bm{x}_{i}$ and compute the difference in the loss of $\bm{x}_{j}$ before and after these gradient updates. We plot these pairwise loss recoveries in Figure 8(b). Results confirm that the amount of recovery on document $\bm{x}_{j}$ is highest when the model checkpoint is taken from roughly $b$ documents before or after document $\bm{x}_{j}$ in cyclic training and then fine-tuned on a proximal document $\bm{x}_{i}$ in the sequence. The fact that this pairwise loss recovery matrix is roughly symmetric also suggests that the anticipatory recovery phenomenon approximately exhibits task symmetry: gradient updates on document $\bm{x}_{t}$ decrease the loss for document $\bm{x}_{t+k}$ for small integers $k$, and vice versa. We provide additional visualizations for $T=50,100,200$ in Appendix C and a more detailed description for this phenomenon.
Figure 9: Top three PCA components of last layer weights in the first three epochs.
(a) $f_{i}(\bm{w})=\bm{w}$
(b) $f_{i}(\bm{w})=\bm{y}_{i}-\bm{w}$
@@ 4.2 Temporal Structure of Model Weights
We explore the structure of model weights along the optimization trajectory of cyclic training. We flatten and concatenate the model weight vectors after fine-tuning on each document. However, the cosine similarities between these raw model weight vectors are all very close to 1 without obvious structure, due to the proximity of model weights along the same optimization trajectory and numerical instability in huge weight vectors. To resolve these issues, we instead explore the structure of “weight residuals.” We compute the weight residuals by subtracting the average of weights in a window of length $T$ centered at the current document from the current weight, i.e. $w_{\text{res}}(t)=w(t)-\frac{1}{T}\sum_{n=t-T/2}^{t+T/2}w(n)$. This removes the shared components along the optimization trajectory and allows us to focus on the model weight updates for each document. Figure 8(c) visualizes a heat map of the cosine similarity between each pair of weight residuals from the second epoch to the fourth epoch. The visualization shows a cyclic structure in the weight residuals, as equidistant bright stripes that align with the training epochs. Furthermore, each stripe spans several documents, suggesting the similarity of weight residuals for proximal documents.
In addition to cosine similarities between weights, we explore visualizing the weights in a lower-dimensional space with Principle Component Analysis (PCA). We compute the top three PCs of the flattened last-layer weight vector (the output word embedding layer) for the Pythia-1B model, and plot its trajectory in Figure 9. The plot exhibits a clear helical structure that gradually converges. We believe this is highly relevant to anticipatory recovery: right before revisiting a task, the projected model weights in the helix move closer to the point corresponding to the previous appearance of that task, leading to anticipatory recovery on the loss of that task. As we go on with cyclic training, the model also exhibits less forgetting and gradually converges to a solution that achieves low loss on all tasks. It is important to note that the helical structure of the weight trajectory is not an obviously necessary consequence of the cyclical training. Cyclical training could be expected to yield a repeating pattern, but the facts that the tasks come to be organized in a circle that respects their ordering and that the trajectory goes through one full revolution per epoch (rather than some other arc length) are nontrivial and seem to be essential for anticipatory recovery.
@@ 4.4 Computational Toy Model
To further understand the essential mechanism that yields anticipatory recovery, we design a minimalist “toy” simulation experiment. In this toy simulation, each task (formerly, document) $i\in\{1,\cdots,T\}$ is described by a single data point, $\bm{x}_{1},\cdots,\bm{x}_{T}\in\mathbb{R}^{N}$. We assume a learnable linear embedding $\bm{P}\in\mathbb{R}^{M\times N}$ that projects each $\bm{x}_{i}$ into an $M$-dimensional embedding space. We also assume a learnable vector $w$ and task-specific mappings $f_{i}$, where $f_{i}(w)$ is the target for task $i$ in the same embedding space. We require each $f_{i}$ to be invertible as a simplifying assumption.
We define the loss for task $i$ as $\ell_{i}(\bm{P},\bm{w})=\frac{1}{2}\lVert\bm{P}\bm{x}_{i}-f_{i}(\bm{w})\rVert^{2}_{2}$. Just as when training a deep net, we assume here that representation learning occurs slowly, and that one training step for task $i$ involves a single gradient update of $\bm{P}$ with step size $\alpha$:
@@
As in our main experiments, we sequentially optimize each $\ell_{i}$ as we iterate through the sequence of tasks. In each training step, we first update $\bm{P}$ and then solve for $\bm{w}$ given the update to $\bm{P}$. Updating $\bm{P}$ approximately reduces the distance between $\bm{P}\bm{x}_{i+1}$ and $f_{i+1}(f_{i}^{-1}(\bm{P}\bm{x}_{i}))$, which entails reducing the upper bound of $||f_{i+1}^{-1}(\bm{P}\bm{x}_{i+1})-f_{i}^{-1}(\bm{P}\bm{x}_{i})||$, assuming that each $f_{i}^{-1}$ is Lipschitz continuous. As a result of this optimization objective, the model will evolve along the optimization trajectory such that the $f_{i}^{-1}(\bm{P}\bm{x}_{i})$ for all tasks $i$ gradually form a circular pattern. This gives an intuitive explanation on the anticipatory recovery phenomenon, since updaing $\bm{w}$ according to equation 2 will also bring it closer to $f_{i+1}^{-1}(\bm{P}\bm{x}_{i+1})$, thus reducing the loss on task $i+1$ and exhibits anticipatory recovery.
@@
There are two potential reasons large overparameterized networks might produce the anticipatory recovery in a way analogous to the toy simulation. First, for larger networks, it is more likely that the network can develop task-specific parameters that quickly adapt to and memorize new input data, corresponding to Equation 2. And when the fast memorization is achieved, the gradient descent dynamics of the slow weights push the representations of the two adjacent tasks ($P\bm{x}_{i}$ and $P\bm{x}_{i+1}$) closer when $f$ is an identity function, according to Equation 1. This effect can be seen in earlier LLM experiments (Figure 2), where larger models achieve significantly lower losses within a few gradient update steps. Second, larger networks have more learning capacity to map the features of two adjacent tasks closer. In our linear projection model, anticipatory recovery keeps growing over many epochs, whereas the anticipatory effect is already at the strongest within 2 or 3 epochs in LLM experiments. Moreover, all data points are randomly generated in the toy model, which makes it easier to separate and map their representations according to a temporal structure than real-world data. In contrast, real-world data could require more representation capacity since data points are noisy and correlated.
@@ 5 Related Work
In this section we discuss the most relevant prior works to this paper. Please refer to Appendix D for additional related work.
@@
Prior theoretical works have studied convergence rates, under various assumptions, for the training setup where the data points are shuffled only once and that order is reused for all epochs (ahn2020sgd, ; gurbuzbalaban2019convergence, ; mishchenko2020random, ; safran2020good, ). On the empirical side, xu2022stochastic found that shuffling the data only once in the beginning can achieve a convergence rate comparable to shuffling every epoch. The training setup is equivalent to our cyclic training setup, but our research examines the loss on each task throughout the training cycle and discovers the anticipatory recovery effect. We also extend it to multiple gradient update steps on each data point.
@@
Online learning deals with the setting where the tasks come from an online sequential stream. One of the simplest algorithms in online learning is follow-the-leader (hannan1957approximation, ), which stores all previous data from the stream and minimizes the total loss. It has strong performance guarantees but is computationally very expensive, and it also might not be feasible to store all the past data. Many subsequent works have developed cheaper algorithms under different assumptions (zinkevich2003online, ; cesa2006prediction, ; shalev2012online, ). Many recent works also explore the connection between online learning and meta-learning or continual learning (denevi2019learning, ; finn2019online, ; denevi2019online, ; javed2019meta, ; fini2020online, ; ren2020wandering, ; wang2021wanderlust, ). The cyclic training setting that we explore in this research can be considered as a special case of the online learning setting where the data stream has a cyclic repetition structure. We employ multiple steps of online gradient descent (biehl1995learning, ) on each document from the stream and study the training dynamics of over-parameterized neural networks.
@@
When transitioning between tasks sequentially, neural networks often experience “catastrophic interference” (mccloskey1989catastrophic, ), marked by a significant drop in performance on previously learned tasks. Numerous algorithms have been proposed to mitigate catastrophic forgetting, focusing on general approaches including parameter regularization (kirkpatrick2017overcoming, ; zenke2017continual, ; aljundi2018memory, ), data replay (rebuffi2017icarl, ; rolnick2019experience, ; chaudhry2019tiny, ), knowledge distillation (hinton2015distilling, ; li2017learning, ; buzzega2020dark, ; madaan2023heterogeneous, ), and architectural isolation and expansion (yoon2017lifelong, ; serra2018overcoming, ; gurbuz2022nispa, ; kang2022forget, ). Our work extends interleaved training (mayo2023multitask, ) to a larger number of tasks, specifically investigating the emergent anticipatory recovery phenomenon in cyclic training. This finding adds to the above literature by demonstrating a new mechanism by which large networks can avoid or recover from catastrophic interference.
Continual Learning
@@
Continual learning chen2018lifelong ; madotto2021continual ; qin2022LFPT5 ; razdaibiedina2023progressive addresses a simplified setup where a model sequentially learns a set of tasks without revision. Recently, there have been debates over the practicality of continual learning setups. Studies like davidson2020sequential have shown that as networks learn more tasks, they improve in learning speed and reduce forgetting. In large models, studies suggest that pre-trained vision classifiers can undertake continual learning with ease, by either freezing or fine-tuning representations (janson2022simple, ; lee2023pre, ; fini2022self, ). In the language domain, research also suggests that LLMs exhibit emerging continual learning capabilities (scialom2022fine, ; ke2022continual, ). Nevertheless, it is uncommon in real environments for tasks to occur only once yet for an agent to need to retain them. Unlike prior literature on continual learning, our research uniquely focuses on sequential learning environments with cyclic repetition.
6 Discussion and Limitations